OpenStudy (anonymous):

A ranger spots a forest fire while on a 45-meter observation tower. The angle of depression from the tower to the fire is 12°. To the nearest meter, how far is the fire from the base of the tower? 212 meters 10 meters 216 meters 71 meters

OpenStudy (campbell_st):

 tan(12) = 45/x x = 45/tan(12)